Catherine Caracciolo – Reflection Paper

Catherine, one of the features of the Montserrat program is that it tries to create a learning community by engaging students both inside and outside the classroom. This means that a few required common events outside of class time are a “part of the deal.” After all, the idea of having all of the students in each cluster live together in one residence is to encourage them to continue the discussion of experiences they have shared beyond the classroom.

Within each cluster (like the Natural World cluster our seminar is a part of) there are admittedly a pretty broad range of different courses, and some common events are inevitably going to be much more directly connected to some courses than others. Since you say that you were interested by the Zetterstrom talk and that you learned some things you had not known about your native county, is there really a problem that the topic was not directly related to the mathematics we are studying in our particular seminar at the moment? You might also be interested to know that we will be directly addressing one aspect of the relation of mathematics to the study of nature (not trees specifically, but the physical world and its contents more generally) later in the semester when we discuss geometry “after Euclid.”

But the main point is that, even if you have not discovered this yet, I think you will find that some of the more interesting and important experiences in life can come from unplanned encounters that don’t fit into preconceived ideas you might have for how things “should work.” I’m not claiming that this talk was necessarily one of those important experiences, by the way! I’m just saying that a little more openness to different things might lead to surprisingly valuable results on occasion.

The main issue I have with this essay is that your writing here seems overly hasty and careless. I don’t have any problem (as I hope you have seen so far) if students express viewpoints that are opposed to the opinions I happen to hold myself. By part of my job is to get you to express the opinions you want to get across in the clearest and most convincing way. This essay does not do that very well because of the many small slips and unclear spots. Look over the corrections and see me if you have any questions.

Grade: Content: 9/10; Writing: 7/10.

Xuan Ha Nghiem – Reflection Paper

Some comments

1. I think that the fearlessness of the Galapagos animals is mainly due to the fact that there are no large land predators in that limited ecosystem. When humans arrived, the native animals literally did not know that they might need to be afraid of and flee from humans.

2. The points on your second page are pretty much a repetition of things you have said before in almost the same way. So I think the essay could have been cut down somewhat in length.

3. I have marked some corrections and suggestions for rewordings. Please see me if you have any questions about those.

Grade: Content: 9/10; Writing: 8/10.

Kaitlin O’Hagan–Reflection Paper

Apart from the fact that the last few sentences are repetitive of things you have said before, this is a good essay on John Hess’s talk on “Darwin’s Tapestry” and the animals of the Galapagos Islands.

I have a comment for you to think about, though. In saying animals have “personalities” (which is, I agree, something that Dr. Hess said himself in the talk), are we maybe reading too much of our own human attitudes into the activities of creatures that are maybe not so much like us after all? This is a serious concern in scientific research about animal behavior, by the way. “Anthropomorphism” is the attribution of human qualities, or human thought patterns, or human motivations to animals, and it is something most scientists try to avoid as much as possible. If you think about it, that might make sense since we might be drawing incorrect conclusions if we fall back on that too much.

Of course it is probably true that Dr. Hess was trying to do something different in his talk. One way to motivate people to care about the environment is to make them feel sympathetic toward the animals and plants. And giving the animals endearing “personalities” is certainly one way to do that.

Grade: Content: 10/10; Writing 9/10.

Peter Zona – Reflection Paper

Your essay on Tom Zetterstrom’s talk is well-written and shows some thought about what he was trying to say.

The main comment I have concerns something you said in your first paragraph. Namely, isn’t it somewhat self-centered for humans to say that trees have a “purpose” in Earth’s ecosystem? I think that biologists, for instance, would shy away from saying that any living thing has a “purpose.” The fact that plants have evolved to take up CO_{2} and give off O_{2} while animals do the reverse is just how it is. And while the system works well for both plants and animals as long as the system is in balance, they would say that neither plants nor animals were designed with that purpose in mind. Of course, there are other ways of looking at things too. Just be sure you appreciate the distinction here!

Grade: Content: 9/10; Writing: 10/10 Kylee Sullivan – Reflection Paper

Oops – most of Tom Zetterstrom’s talk dealt with American elm trees, not oak trees. They are a completely different species. Small slips like this might not seem that important, but they can really sabotage the effect you want your writing to have(!) So please be more careful.

Comments

1. The “shade” issue is a very important one from the point of view of humans who use trees to try to control their environments. One of the main features of elm trees was that they did provide a lot of shade. That meant that yards and houses shaded by elms stayed cooler in the summer. Don’t forget that in the time that elm trees were flourishing in the U.S. there was virtually no air conditioning around. So this was actually a very important reason why elms were so commonly planted in U.S. cities and towns.

2. I don’t remember Zetterstrom criticizing the lack of variety of trees on the Holy Cross campus. As I recall, his main point there was that our trees are sort of “museum pieces” rather than being integrated into all facets of the campus (for instance, the lack of trees in the expanses of the Hogan Campus Center parking lots). I think he also made the point that we don’t use the trees we have as well as we could as a teaching resource (the idea of a “teaching arboretum.”)

Grade: Content: 8.5/10 ; Writing: 10/10.

Tom Santa Maria – Reflection Paper

I’m not sure I follow the final sentence in your first paragraph. It seems to be going in two different directions at the same time. One possibility is that you really should start with “Though I can absolutely and honestly say ... ” That would be the way to say unequivocally that the talk did nothing for you. Since you’re not really saying that, I think you really meant “Despite my attitude, and although I can absolutely and honestly say that I did not emerge from Seelos Theater a changed man, one message ... ” The point is that the “Despite my attitude ” signals that there was some change from what you thought before, and that in some way, contrary to your expectations, the talk did change your mind. So this should be clearer.

Grade: Content: 10/10; Writing 9.5/10.

Andrew Marzo – Reflection Paper

This is a very good essay about the main points that Tom Zetterstrom was trying to get across in his talk. There are just a couple of technical issues in the writing. First, when you are referring to the species of tree, elm is not capitalized. Second, the next to last sentence needs some work. I’m not sure how “distribute” fits with what you are trying to say. It seemed to me that “upset” or “disrupt” would be closer to what I think you mean. If on the other hand you are referring to the power humans have to spread particular plants and animals (or diseases) around then say something like “spread a disease,” not “distribute an ecosystem.”

Grade: Content: 10/10; Writing 9/10.

Max Pettinelli – Reaction Paper

I’m very glad to hear that the photography exhibit caught your interest. I was really struck by the series of photographs of the elm tree across the yearly cycle (in the four seasons) as well. They are a reminder how tenacious, and at the same time how fragile, life is. Your short essay is also very well-written. Good work!

Grade: Content 10/10; Writing: 10/10.

Lucia Bostrom – Reaction Paper

Your short essay is very well-written and I’m glad the talk gave you the opportunity to think about these things. The one comment I have is that nature is not quite as defenceless against humans as you seem to indicate toward the end of the essay. Think of the way disease germs can acquire immunity to the drugs that we use to kill them, or the way bedbugs (ick!) seem to be making a major comeback in cities across the U.S. right now! In my opinion, the more we try to live with and in the natural world rather than trying to tailor it exclusively to our liking, the happier we will be in the long run as a species.

Grade: Content: 10/10; Writing: 10/10.

Astrid Ludwig – Reaction Paper

This is a good essay giving a reaction to Dr. John Hess’s presentation on “Darwin’s Tapestry” and the animals of the Galapagos Islands. I have a couple of comments, though. The first deals with the way you say that the animals in that ecosystem have to “work for each other.” It was not entirely clear to me what you meant by that. If you mean that they are interdependent, then I would certainly agree. But I think it would be a big “stretch” to say that they have any conscious collaboration to make their lives possible.

The second comment is that your statements about how isolated we are from the natural world certainly apply to how we live in the U.S. at this point in history. But if you think about it, this is actually a very different sort of life from the lives most humans have lived for most of our time on the earth. Almost everyone lived in close proximity with domesticated animals like cows, goats, sheep, pigs, etc., for instance, before the rise of urban life. And many people around the world still live that way today. Moreover, people also had to cope with wild animals much more frequently and in much more serious ways than we do now. Consider the situation of settlers in newly-opened frontier regions of the U.S. in the 1800’s, for instance. Those settlers probably had to deal with wild predators like wolves killing their farm animals, or wild birds and deer eating some of their garden crops, or bears foraging for food around their homes on a pretty regular basis. My point is that you need to understand how different things are now than they were for most of our (humans’) history.

Grade: Content: 9.5/10; Writing: 9.5/10.

Dan Finn – Reflection Paper

Your essay in reaction to Tom Zetterstrom’s talk is well-written for the most part. I hope you will follow up on your sense that we should all find ways to participate in the work of preserving the natural world. If you are interested in pursuing academic work along these lines you might consider the Environmental Studies concentration. There are also ways to get involved in service activities related to the environment if you are inclined that way.

I also have a couple of technical comments for you. First, in the phrase “Dutch elm disease” only the first word is capitalized (since it is the adjective referring to the country). Second, most people think that the disease was introduced into the U.S. in a shipment of wood from the Netherlands (not France as you said).

Grade: Content 10/10; Writing: 9.5/10 Muhammad Juvaid – Reaction Paper

Your essay in reaction to Tom Zetterstrom’s talk is well-written for the most part. There is one sentence fragment, though, in the first paragraph. I would say that you should try to avoid these in formal writing (even though I know that we tend sometimes to add a point without including a subject and a verb when we are speaking). It was was also not completely clear to me what you meant to say in that sentence. Did you mean to say that elm trees have a lot of historical importance in parts of the U.S. (such as Massachusetts, where it is our state tree)? Another smaller point is that the name “elm” is not capitalized when you are just referring to the species of tree.

I hope you will follow up on your sense that we should all find ways to participate in the work of preserving the natural world. If you are interested in pursuing academic work along these lines you might consider the Environmental Studies concentration. There are also ways to get involved in service activities related to the environment if you are inclined that way.

Grade: Content: 9.5/10; Writing: 9/10.

John Macomber – Reaction Paper

I agree that the natural world can be a source of calm and rejuvenation for human beings. I know that you are not exactly saying what I am going to mention next. Some humans tend to think that because nature can have a calming, rejuvenating effect on them, that is somehow its purpose or reason for being there. Of course, it is arrogant to think that way. Humans are part of nature and we cannot control everything that happens in nature to make it into a perfect parkland for our own enjoyment and refreshment.

On the technical side, there is a sentence fragment in the first paragraph. I would say that you should try to avoid these in formal writing (even though I know that we tend sometimes to add a point like this without including a subject and a verb when we are speaking). There are also a few smaller slips that I have marked with corrections. If you have any questions about these, please see me.

Grade: Content: 10/10 ; Writing: 8.5/10.

Kyle Kadlick – Reaction Paper

The main comment I have concerns something you said toward the beginning of the essay. Isn’t it somewhat self-centered for humans to say that trees have a purpose in Earth’s ecosystem? I think that biologists, for instance, would shy away from saying that any living thing has a “purpose.” The fact that plants have evolved to take up CO_{2} and give off O_{2}, while animals do the reverse is just how it is. While the system works well for both plants and animals while it is in balance, biologists would deny that either plants or animals were designed with that purpose in mind, or only designed in the most general, impersonal way by the forces of evolution.

The sentence I marked on the paper could use some polishing too. Your phrase “proclaim about” is not correct since “proclaim” always takes a direct object. You can proclaim a holiday or proclaim a new constitution, but you cannot “proclaim about trees.”

If you want to follow up on your sense that we should all find ways to participate in the work of preserving the natural world, there are several ways to do that at Holy Cross. If you are interested in pursuing academic work along these lines you might consider the Environmental Studies concentration. There are also ways to get involved in service activities related to the environment if you are inclined that way.

Grade: Content: 9/10; Writing: 9/10.

Sarah Cahn – Reaction Paper

An arboretum is a collection of trees maintained together by an organization of people. Tom Zetterstrom’s discussion of the fact that Holy Cross’s campus arboretum is not being used as fully as it could be was a good point to include here. He suggested that the campus arboretum could be used for teaching in many ways that are not happening now (at least not in any organized way). He meant by this that all of our students could be educated more about the trees on campus as a way of introducing them to aspects of the natural world. At the same time, the trees might be used more fully as a resource for projects in biology and other courses so that they would become a resource for educational activities on campus. It would have been good to expand this point more fully in your essay, since just saying “teaching arboretums” does not really get across all the interesting ways the College’s collection of trees might be used.

Grade: Content: 9/10; Writing: 10/10.

Matt Kasuba – Reaction Paper

If you want to follow up on your sense that we should all find ways to participate in the work of preserving the natural world, there are several ways to do that at Holy Cross. If you are interested in pursuing academic work along these lines you might consider the Environmental Studies concentration. There are also ways to get involved in service activities related to the environment if you are inclined that way.

Grade: Content 10/10; Writing: 9.5/10.

Kevin Sutman – Reaction Paper

One interesting aspect of Tom Zetterstrom’s talk was his idea that Holy Cross’s campus arboretum is not being used as fully as it could be suggested that the campus arboretum could be used for teaching. He meant by this that all of our students could be educated more about the trees on campus as a way of introducing them to aspects of the natural world. At the same time, the trees might be used more fully as a resource for projects in biology and other courses so that they would become a resource for educational activities on campus.

If you want to follow up on your sense that we should all find ways to participate in the work of preserving the natural world, there are several ways to do that at Holy Cross. If you are interested in pursuing academic work along these lines you might consider the Environmental Studies concentration. There are also ways to get involved in service activities related to the environment if you are inclined that way.

Grade: Content 10/10; Writing 10/10.